Drive the following formulas;

d/dx [xvYv(x)]=xvYv-1(x)


1
Expert's answer
2021-03-09T06:00:49-0500

Taking LHS


"\dfrac{d}{dx}[x^vY_v(x)]"


Now let "m = Y_v(x)" , "n= x^v"


Now using the Leibnitz theorem


"\dfrac{d}{dx}[x^vY_v(x)]"


="^1C_0(1!\times x^v \times Y_{v-1}(x))"


= "x^vY_{v-1}(x)"